# Successive randomized compression: A randomized algorithm for the compressed MPO–MPS product

### v1

@@ 1 Introduction
    MPSs are used to represent states of one-dimensional quantum systems such as chains of interacting spin-$S$ particles. In other applications, they are used to represent vectors or functions. MPSs are also called tensor trains, particularly in the applied mathematics literature.
@@
[Uncaptioned image][Uncaptioned image].[Uncaptioned image][Uncaptioned $\end{gathered}.$ (1.1)
@@
Given the importance of the compressed MPO–MPS multiplication problem, many algorithms have been designed for the task; see section 4 for a description of existing methods. A comparison of these methods is shown in Figure 1, which shows the relative error versus the runtime of several methods for computing the product of a random MPO and a random MPS. We set both bond dimensions to $D=\chi=50$, use $n=100$ sites, and consider a physical dimension $d=2$. Each tensor is populated with uniformly random entries between $[\alpha,1]$ for $\alpha=-0.5$. We use complex data types for this experiment. The parameter $\alpha\in[-1,1]$ sets the difficulty of the problem, with higher $\alpha$ making the problem easier [31]; we choose $\alpha=-0.5$ for an intermediate level of difficulty. This is a non-physical example purely intended to demonstrate the performance and accuracy characteristics of the existing methods.
Figure 1: (MPO–MPS contraction algorithm comparison). Comparison of relative error versus computation time of several algorithms for the MPO–MPS product of $n=100$ tensor networks with bond dimension $D=\chi=50$. The maximum bond dimension $\overline{\chi}$ was varied between $5$ and $100$, with the compressibility difficulty parameter $\alpha$ set to $-0.5$. The fitting (variational) method was run for a single left-to-right sweep; more sweeps led to higher computational cost at no appreciable gain in accuracy. Randomized C-T-C is the contract-then-compress method with randomized MPS rounding (see section 4.1.2). Each data point represents the mean of five independent runs.
Figure 1 demonstrates limitations of the existing MPO–MPS algorithms. The contract-then-compress algorithm (also called the “naïve” algorithm), its randomized variant, and the density matrix algorithm are all accurate but slow (at least on this problem with $D,\chi\gg 1$). The zip-up method, by contrast, is fast but less accurate. On this example, the fitting (i.e., variational) method is accurate but slower than the zip-up method. As we will see in section 4.2.1, the fitting method can require many sweeps to converge (making it significantly slower), or it may fail to converge at all.
@@
  • Accurate. For a given bond dimension, the output $|\eta\rangle\approx H|\psi\rangle$ of the SRC algorithm achieves accuracy comparable to the contract-then-compress method, which is known to achieve near-optimal truncation error [32, Cor. 2.4]. In particular, the accuracy of SRC is often significantly better than the zip-up algorithm.
@@
We believe this suite of benefits make the SRC method a compelling candidate for deployment in applications. Indeed, in Figure 1, SRC achieves higher accuracy in less time than other randomized MPO–MPS methods.
@@ 1.1 Outline for paper
After introducing background material on randomized matrix approximation and Khatri–Rao products in section 2, we present our algorithm in section 3. Description of and comparison with existing MPO–MPS product methods appears in section 4, and an application to unitary time evolution using the TDVP1-GSE algorithm appears in section 5.
@@ 1.2 Notation
The Frobenius norm of a matrix or tensor is $\left\|\cdot\right\|_{\rm F}$. We denote the Hermitian adjoint of a matrix by . The symbol $\otimes$ refers to the Kronecker product of vectors or matrices, while $\odot$ denotes the Khatri–Rao product of matrices (see section 2.2). We express tensors $|\psi\rangle\in\mathbb{C}^{d\times\cdots\times d}$ using bra–ket notation. The constituent tensors of an MPS $|\psi\rangle$ are denoted $\psi^{(1)},\ldots,\psi^{(n)}$ (similarly, $H^{(i)}$ for an MPO $H$).
@@ 2 Background
This section reviews background material for our randomized method for the compressed MPO–MPS multiplication algorithm, including randomized low-rank approximation (section 2.1) and the Khatri–Rao product (section 2.2).
@@ 2.1 Randomized QB approximation
Randomized low-rank approximation algorithms [33, 34, 35, 36] are a class of fast, robust methods for approximating a large matrix $A\in\mathbb{C}^{M\times N}$ as a product $A\approx QB$ of thin matrix $Q\in\mathbb{C}^{M\times p}$ times a wide matrix $B\in\mathbb{C}^{p\times N}$. In this section, we review the simplest of such methods, randomized QB approximation. In the next section (section 3), we will show how the randomized QB approximation can be used to give a fast algorithm for the MPO–MPS product.
@@
The randomized QB approximation $\widehat{A}=QB$ is the same as the approximation produced by the randomized SVD algorithm [33], but presented in the form $\widehat{A}=QB$ rather than as an SVD. When the randomized QB approximation is exact ($A=\widehat{A}$), we call the expression $A=QB$ a (randomized) QB decomposition.
The accuracy of randomized QB approximation can be understood using the following result [33, Thm 10.5] (see also [37, Fact A.2]):
@@
The conclusion of this result is that randomized QB approximation exactly recovers genuinely low-rank matrices and, with modest oversampling such as $p=r+5$ or $p=2r$, produces an approximation that is comparable with the best rank-$r$ approximation. See [33, §10], [35, §8], and the references therein for many more analyses of randomized QB approximation. While most of the analysis of randomized QB approximation focuses on the Gaussian case, the method usually works quite well in practice for a much richer class of random matrices $\Omega$ [38, §11.5.2].
@@ 2.2 The Khatri–Rao product
To perform QB approximation efficiently for matrix unfoldings of tensor networks, we can use a tensor-structured random matrix $\Omega$. There are many different types of tensor-structured random matrices that have been proposed for linear algebraic computations; see [36, Ch. 7] for a survey. In this paper, we will use the simplest example, a Khatri–Rao product of random matrices.
The Khatri–Rao product [39, §2.6] is defined as follows. Let $\Omega^{(1)},\ldots,\Omega^{(n)}\in\mathbb{C}^{d\times p}$ be matrices, where each matrix $\Omega^{(j)}$ has columns
@@
To apply randomized QB approximation to MPO–MPS products in this paper, we will define $\Omega\coloneqq\Omega^{(1)}\odot\cdots\odot\Omega^{(n)}$ to be a Khatri–Rao product of random matrices $\Omega^{(1)},\ldots,\Omega^{(n)}$ with standard normal entries, all mutually independent.
@@
We emphasize in this diagram that the white circle at the top of this diagram indicates multiplication of the corresponding indices entrywise. See [40, §1.2 & §3.7] for more details on denoting Khatri–Rao products using tensor diagrams.
@@
Appendix A contains a proof of Theorem 2. We expect that this result appears in the literature, but we were unable to locate a citation. Unfortunately, a quantitative error bound of the form 2.1 is not available when $\Omega$ is a Khatri–Rao product.
@@ 3 Successive randomized compression
In practice, the goal is typically to compute an approximate compression $|\eta\rangle\approx H|\psi\rangle$ and, often, to determine $\overline{\chi}$ adaptively to satisfy an error tolerance. The algorithm works without modification for approximate contraction (though a final rounding step can be helpful, see section 3.4). We describe the algorithm in detail in sections 3.1, 3.2 and 3.3, and we analyze its computational cost in section 3.5. Adaptive determination of bond dimension is deferred to appendix C.
@@ 3.1 Step 1: The last site
Begin by viewing $H|\psi\rangle$ as a $d\times d^{n-1}$ matrix, with the last visible index as the rows and the first $n-1$ indices as the columns:
@@
$(H|\psi\rangle)\cdot\Omega^{(1:n-1)}$ =[Uncaptioned image].absent[Uncaptioned $\end{gathered}\,.$
@@
$(H|\psi\rangle)\cdot\Omega^{(1:n-1)}$ =[Uncaptioned image].absent[Uncaptioned $\end{gathered}\,.$
@@
$(H|\psi\rangle)\cdot\Omega^{(1:n-1)}$ =[Uncaptioned image].absent[Uncaptioned $\end{gathered}.$
@@
$(H|\psi\rangle)\cdot\Omega^{(1:n-1)}$ =[Uncaptioned image]=[Uncaptioned image].absent[Uncaptioned image][Uncaptioned $\end{gathered}\,.$
@@
Since the right index of $Y^{(n)}$ corresponds to the rows, the orthogonal “$\mathsf{Q}$” factor $\eta^{(n)}$ occurs to the right of the triangular “$\mathsf{R}$” factor $R^{(n)}$ in the tensor diagram 3.1, but the factorization is written $Y^{(n)}=\eta^{(n)}\cdot R^{(n)}$ in matrix notation. The tensor $\eta^{(n)}$ will form the final site of the output $|\eta\rangle=H|\psi\rangle$.
@@
Remember that we are assuming, for now, that $H|\psi\rangle$ is exactly compressible to an MPS of bond dimension $\overline{\chi}$. Thus, by Theorem 2, the randomized QB decomposition is exact with probability one.
In the next step of the algorithm, we will produce site $n-1$ of $H|\psi\rangle$ by applying randomized QB decomposition to $B^{(n-1)}$. In preparation for this step, we contract $(\eta^{(n)})^{\dagger}$ into the MPO and MPS sites below, obtaining
@@ 3.2 Step 2: Second-to-last site
We now compute a randomized QB decomposition of $B^{(n-1)}$. We treat $B^{(n-1)}$ as a $\overline{\chi}d\times d^{n-2}$ matrix, with the last two exposed indices as rows and the first $n-2$ visible indices as columns:
The key step which allows our algorithm to be computationally efficient is reuse of the same random matrices $\Omega^{(1)},\ldots,\Omega^{(n-2)}$ used in the first step of the algorithm. Specifically, we form the Khatri–Rao product $\Omega^{(1:n-2)}\coloneqq\Omega^{(1)}\odot\cdots\odot\Omega^{(n-2)}$ and apply it to $B^{(n-1)}$, resulting in the following tensor network:
@@
Observe that $B^{(n-2)}$ now has $n-2$ exposed indices of dimension $d$ and $1$ exposed index of dimension $\overline{\chi}$; the total number of exposed indices, $n-1$, has been decreased by one.
@@ 3.4 Optional step: Oversampling and final round
In sections 3.1, 3.2 and 3.3, we introduced the SRC algorithm under the assumption that the product $H|\psi\rangle$ is exactly representable as an MPS of known bond dimension $\overline{\chi}$. However, the algorithm works without modification for computing a compressed approximation $H|\psi\rangle$ as an MPS with specified bond dimension $\overline{\chi}$.
@@ 3.5 Summary: Pseudocode and time complexity
1:MPO $H$, MPS $|\psi\rangle$, and output bond dimension $\overline{\chi}$ (or tolerance $\tau$, see section C.2)
@@
14:[Optional: Compute error estimate and increase $\overline{\chi}$ if needed; see section C.2]
@@
21:     [Optional: Compute error estimate and increase $\overline{\chi}$ if needed; see section C.2]
@@
This result follows directly from Theorem 2, though the fact that we use a common set of random matrices $\Omega^{(1)},\ldots,\Omega^{(n-1)}$ across all steps of the algorithms makes the result not entirely trivial. Appendix B contains a proof of this result.
When $H|\psi\rangle$ is not exactly equal to an MPS of bond dimension $\overline{\chi}$, SRC returns a compressed approximation $|\eta\rangle\approx H|\psi\rangle$ to the product. We present pseudocode for SRC in Algorithm 1. See appendix C for a discussion of determining the output bond dimension $\overline{\chi}$ adaptively to meet a tolerance.
@@
Using either the full runtime 3.3a or simplified runtime 3.3b, SRC is as fast as or faster than all existing MPO–MPS multiplication algorithms; see section 4 for a description of other MPO–MPS multiplication algorithms and a runtime comparison with SRC.

### v2

# Successive randomized compression: A randomized algorithm for the compressed MPO-MPS product

@@ 1 Introduction
    MPSs are used to represent states of one-dimensional quantum systems such as chains of interacting spin-$S$ particles. In other applications, they are used to represent vectors or functions. MPSs are also called tensor trains (TTs), particularly in the applied mathematics literature.
@@
[…]
@@
Given the importance of the compressed MPO–MPS multiplication problem, many algorithms have been designed for the task; see section˜4 for a description of existing methods. A comparison of these methods appears in Figure˜1, which shows the relative error $\|H|\psi\rangle-|\eta\rangle\|_{\rm F}/\|H|\psi\rangle\|_{\rm F}$ versus the runtime of several methods for computing the product of a random MPO and a random MPS. The Frobenius norm $\left\|\cdot\right\|_{\rm F}$ is defined as the square root of the sum of squared entries in any tensor. We set both bond dimensions to $D=\chi=50$, use $n=100$ sites, and consider a physical dimension $d=2$. Each tensor is populated with uniformly random entries in the range $\alpha,1]$ for $\alpha=-0.5$. The data for this problem is real, but we store it using complex data types (which is the required data format for most quantum tensor network problems). The parameter $\alpha\in[-1,1]$ sets the difficulty of the problem, with higher $\alpha$ making the problem easier [31, 32, 33]; we choose $\alpha=-0.5$ for an intermediate level of difficulty. This is a non-physical example purely intended to demonstrate the performance and accuracy characteristics of the existing methods.
Figure 1: (MPO–MPS contraction algorithm comparison). Comparison of relative error versus computation time of several algorithms for the MPO–MPS product of $n=100$ tensor networks with bond dimension $D=\chi=50$. The maximum bond dimension $\overline{\chi}$ was varied between 55 and $100$, with the compressibility difficulty parameter $\alpha$ set to 0.5-0.5. The fitting (variational) method was run for a single left-to-right sweep; more sweeps led to higher computational cost at no appreciable gain in accuracy. C-T-C is an abbreviation for “contract-then-compress”. Randomized C-T-C refers to the contract-then-compress method with randomized MPS rounding (see section˜4.1.2). Each data point represents the mean of five independent runs.
Figure˜1 demonstrates limitations of the existing MPO–MPS algorithms. The contract-then-compress (C-T-C) algorithm (also called the “naïve” algorithm), and the randomized contract-than-compress algorithm, and the density matrix algorithm are all accurate but slow (at least on this problem with $D,\chi\gg 1$). The zip-up method, by contrast, is fast but less accurate. On this example, the fitting (i.e., variational) method is accurate but slower than the zip-up method. As we will see in section˜4.2.1, the fitting method can require many sweeps to converge (making it significantly slower), or it may fail to converge at all.
@@
  • Accurate. For a given bond dimension, the output $|\eta\rangle\approx H|\psi\rangle$ of the SRC algorithm achieves accuracy comparable to the contract-then-compress method, which is known to achieve near-optimal truncation error [34, Cor. 2.4]. In particular, the accuracy of SRC is often significantly better than the zip-up algorithm.
@@
We believe this suite of benefits make the SRC method a compelling candidate for deployment in applications. Indeed, in Figure˜1, SRC achieves higher accuracy in less time than other randomized MPO–MPS methods.
@@ 1.1 Outline for paper
After introducing background material on randomized matrix approximation and Khatri–Rao products in section˜2, we present our algorithm in section˜3. Description of and comparison with existing MPO–MPS product methods appears in section˜4, and an application to unitary time evolution using the TDVP1-GSE algorithm appears in section˜5.
@@ 1.2 Notation
The Frobenius norm of a matrix or tensor is $\left\|\cdot\right\|_{\rm F}$. We denote the Hermitian adjoint of a matrix by . The symbol $\otimes$ refers to the Kronecker product of vectors or matrices, while $\odot$ denotes the Khatri–Rao product of matrices (see section˜2.2). We express tensors $|\psi\rangle\in\mathbb{C}^{d\times\cdots\times d}$ using bra–ket notation. The constituent tensors of an MPS $|\psi\rangle$ are denoted $\psi^{(1)},\ldots,\psi^{(n)}$ (similarly, $H^{(i)}$ for an MPO $H$).
@@ 2 Background
This section reviews background material for our randomized method for the compressed MPO–MPS multiplication algorithm, including randomized low-rank approximation (section˜2.1) and the Khatri–Rao product (section˜2.2).
@@ 2.1 Randomized QB approximation
Randomized low-rank approximation algorithms [35, 36, 37, 38] are a class of fast, robust methods for approximating a large matrix $A\in\mathbb{C}^{M\times N}$ as a product $A\approx QB$ of thin matrix $Q\in\mathbb{C}^{M\times p}$ times a wide matrix $B\in\mathbb{C}^{p\times N}$. In this section, we review the simplest of such methods, randomized QB approximation. In the next section (section˜3), we will show how the randomized QB approximation can be used to give a fast algorithm for the MPO–MPS product.
@@
The randomized QB approximation $\widehat{A}=QB$ is the same as the approximation produced by the randomized SVD algorithm [35], but presented in the form $\widehat{A}=QB$ rather than as an SVD. The matrix $Q$ has orthonormal columns, and the matrix $B$ has no special structure. When the randomized QB approximation is exact ($A=\widehat{A}$), we call the expression $A=QB$ a (randomized) QB decomposition.
The accuracy of randomized QB approximation can be understood using the following result [35, Thm 10.5] (see also [39, Fact A.2]):
@@
The conclusion of this result is that randomized QB approximation exactly recovers genuinely low-rank matrices and, with modest oversampling such as $p=r+5$ or $p=2r$, produces an approximation that is comparable with the best rank-$r$ approximation. See [35, §10], [37, §8], and the references therein for many more analyses of randomized QB approximation. While most of the analysis of randomized QB approximation focuses on the Gaussian case, the method usually works quite well in practice for a much richer class of random matrices $\Omega$ (see [40, §11.5.2] and [41, Fig. 1]).
The accuracy of randomized QB approximation can be improved by using a random matrix $\Omega$ that is more aligned with the “important directions” (i.e., right singular vectors) of the matrix $A$. Such a better $\Omega$ can be obtained systematically by applying subspace or Krylov iteration [35, 42, 43, 37] or by using prior knowledge about the matrix [44]. In the context of this work, subspace iteration and Krylov iteration are of prohibitively high cost, and we assume no prior information is available. Therefore, a random $\Omega$, generated independently from the matrix $A$, is the best-available choice.
Randomized QB approximation or the randomized SVD can be used as a replacement for a truncated full SVD in any algorithm, and there are several papers [45, 46, 47, 48] that have proposed using the randomized SVD as a fast alternative to full SVDs in tensor network algorithms. In this paper, we go further by applying the randomized QB approximation directly to an entire tensor network.
@@ 2.2 The Khatri–Rao product
To perform QB approximation efficiently for matrix unfoldings of tensor networks, we can use a tensor-structured random matrix $\Omega$. There are many different types of tensor-structured random matrices that have been proposed for linear algebraic computations; see [38, Ch. 7] for a survey. In this paper, we will use the simplest example, a Khatri–Rao product of random matrices.
The Khatri–Rao product [49, §2.6] is defined as follows. Let $\Omega^{(1)},\ldots,\Omega^{(n)}\in\mathbb{C}^{d\times p}$ be matrices, where each matrix $\Omega^{(j)}$ has columns
@@
To apply randomized QB approximation to MPO–MPS products in this paper, we will define $\Omega\coloneqq\Omega^{(1)}\odot\cdots\odot\Omega^{(n)}$ to be a Khatri–Rao product of random matrices $\Omega^{(1)},\ldots,\Omega^{(n)}$ with standard (real or complex) normal entries, all mutually independent. The resulting Khatri–Rao random matrix $\Omega$ consists of a collection of random product states $\omega_{j}^{(1)}\otimes\omega_{j}^{(2)}\otimes\cdots\otimes\omega_{j}^{(n)}$, stacked columnwise. We can compute the action $A\Omega$ of a matrix $A$ on $\Omega$ by applying $A$ to each product state, one at a time.
@@
We emphasize in this diagram that the white circle at the top of this diagram indicates multiplication of the corresponding indices entrywise. See [50, §1.2 & §3.7] for more details on denoting Khatri–Rao products using tensor diagrams.
@@
Appendix˜A contains a proof of Theorem˜2. We expect that this result appears in the literature, but we were unable to locate a citation. After the initial release of this paper, the authors developed a quantitative error bound for randomized QB approximation, similar to ˜2.1, in collaboration with Raphael Meyer [41, §5].
The random test matrices with Khatri–Rao product structure used in this work may be contrasted with the framework of “tensor train sketching” (e.g., see [51, 14, 52, 53, 54]). In the latter approach, one chooses $\Omega$ to be a random MPS with a single extra exposed index at the right-most site, which has dimension $p$:
The tensor train sketching framework has advantages and drawbacks over the random Khatri–Rao product matrices used in this work. As advantages, the user has more freedom, as they can tune the bond dimensions of this MPS as a way of trading off the accuracy of the algorithm against computational cost. As disadvantages, the computational effort of tensor train sketching is larger than random Khatri–Rao test matrices, and setting the bond dimensions of the tensor train sketching matrix generally requires either problem-dependent tuning or prior knowledge about the problem.
@@ 3 Successive randomized compression
In practice, the goal is typically to compute an approximate compression $|\eta\rangle\approx H|\psi\rangle$ and, often, to determine $\overline{\chi}$ adaptively to satisfy an error tolerance. The algorithm works without modification for approximate contraction (though a final rounding step can be helpful, see section˜3.4). We describe the algorithm in detail in sections˜3.1, 3.2 and 3.3, and we analyze its computational cost in section˜3.5. Adaptive determination of bond dimension is deferred to appendix˜C.
@@ 3.1 Step 1: The last site
Begin by viewing $H|\psi\rangle$ as a $d\times d^{n-1}$ matrix, with the last visible index as the rows and the first n1n-1 indices as the columns:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Since the right index of $Y^{(n)}$ corresponds to the rows, the orthogonal $\mathsf{Q}$” factor $\eta^{(n)}$ occurs to the right of the triangular $\mathsf{R}$” factor $R^{(n)}$ in the tensor diagram ˜3.1, but the factorization is written $Y^{(n)}=\eta^{(n)}\cdot R^{(n)}$ in matrix notation. The tensor $\eta^{(n)}$ will form the final site of the output $|\eta\rangle=H|\psi\rangle$.
@@
Remember that we are assuming, for now, that $H|\psi\rangle$ is exactly compressible to an MPS of bond dimension $\overline{\chi}$. Thus, by Theorem˜2, the randomized QB decomposition is exact with probability one.
In the next step of the algorithm, we will produce site n1n-1 of $H|\psi\rangle$ by applying randomized QB decomposition to $B^{(n-1)}$. In preparation for this step, we contract $(\eta^{(n)})^{\dagger}$ into the MPO and MPS sites below, obtaining
@@ 3.2 Step 2: Second-to-last site
We now compute a randomized QB decomposition of $B^{(n-1)}$. We treat $B^{(n-1)}$ as a $\overline{\chi}d\times d^{n-2}$ matrix, with the last two exposed indices as rows and the first n2n-2 visible indices as columns:
The key step which allows our algorithm to be computationally efficient is reuse of the same random matrices $\Omega^{(1)},\ldots,\Omega^{(n-2)}$ employed in the first step of the algorithm. Specifically, we form the Khatri–Rao product $\Omega^{(1:n-2)}\coloneqq\Omega^{(1)}\odot\cdots\odot\Omega^{(n-2)}$ and apply it to $B^{(n-1)}$, resulting in the following tensor network:
@@
Observe that $B^{(n-2)}$ now has n2n-2 exposed indices of dimension $d$ and $1$ exposed index of dimension $\overline{\chi}$; the total number of exposed indices, n1n-1, has been decreased by one.
@@ 3.4 Optional step: Oversampling and final round
In sections˜3.1, 3.2 and 3.3, we introduced the SRC algorithm under the assumption that the product $H|\psi\rangle$ is exactly representable as an MPS of known bond dimension $\overline{\chi}$. However, the algorithm works without modification for computing a compressed approximation $H|\psi\rangle$ as an MPS with specified bond dimension $\overline{\chi}$.
@@ 3.5 Summary: Pseudocode and time complexity
1:MPO $H$, MPS $|\psi\rangle$, and output bond dimension $\overline{\chi}$ (or tolerance $\tau$, see section˜C.2)
@@
14:[Optional: Compute error estimate and increase $\overline{\chi}$ if needed; see section˜C.2]
@@
21:  [Optional: Compute error estimate and increase $\overline{\chi}$ if needed; see section˜C.2]
@@
This result follows directly from Theorem˜2, though the fact that we use a common set of random matrices $\Omega^{(1)},\ldots,\Omega^{(n-1)}$ across all steps of the algorithms makes the result not entirely trivial. Appendix˜B contains a proof of this result.
When $H|\psi\rangle$ is not exactly equal to an MPS of bond dimension $\overline{\chi}$, SRC returns a compressed approximation $|\eta\rangle\approx H|\psi\rangle$ to the product. We present pseudocode for SRC in Algorithm˜1. See appendix˜C for a discussion of determining the output bond dimension $\overline{\chi}$ adaptively to meet a tolerance.
@@
Using either the full runtime ˜3.3a or simplified runtime ˜3.3b, SRC is as fast as or faster than all existing MPO–MPS multiplication algorithms; see section˜4 for a description of other MPO–MPS multiplication algorithms and a runtime comparison with SRC.
3.6 Extension: Linear combinations of MPO–MPS products
As an extension, this algorithm can also be used to form a compressed representation of a linear combination of MPO–MPS products:
$|\eta\rangle\approx\sum_{i=1}^{t}\alpha H_{i}|\psi_{i}\rangle.$
At each phase of the algorithm, we compute a “local” $Y^{(j)}_{i}$ tensor separately for each term, and compute the global $Y^{(j)}=\sum_{i=1}^{t}\alpha_{i}Y^{(j)}_{i}$ as a linear combination. The global $Y^{(j)}=\eta^{(j)}\cdot R^{(j)}$ tensor is then split using a QR factorization, and we use the same $\eta^{(j)}$ to define a local tensor $B^{(j-1)}_{i}=(\eta^{(j)})^{\dagger}B^{(j)}_{i}$ for each term. We conclude the algorithm by setting $\eta^{(1)}\coloneqq\sum_{i=1}^{t}\alpha_{i}B^{(1)}_{i}$. This randomized algorithm for compressing sums of MPO–MPS products is similar to the algorithm for compressing sums of MPSs in [51].
3.7 Limitation: Physical symmetries
In physical problems, MPOs and MPSs often obey group symmetry relations associated with physical conservation laws (e.g., symmetry under the group $U(1)$ arises from particle-number conservation). When symmetry is present, it can be built into the tensor network representation by requiring that the constituent tensors $\psi^{(1)},\ldots,\psi^{(n)}$ or $H^{(1)},\ldots,H^{(n)}$ have a prescribed block-sparsity structure; see [55, 56] for details.
Unlike several other compressed contraction algorithms, the SRC algorithm presented in this paper is not symmetry-preserving: The intermediate tensors $C^{(i)}$ and $S^{(i)}$ computed by the algorithm do not possess sparsity structure, and the algorithm does not ensure the compressed output $|\eta\rangle$ has the appropriate block-sparsity structure. (The algorithm can, however, exploit sparsity of the constituent tensors by using sparse tensor formats to perform intermediate contractions in the algorithm.) Investigating whether the SRC algorithm can be modified to respect symmetries is a natural subject for future research.